# Localized Orthogonal Decomposition Methods vs. Classical FEM for the Gross-Pitaevskii Equation

### v1

@@ 3 Spatial Discretization by Localized Orthogonal Decomposition
A natural choice for a basis of $V_{\text{\tiny LOD}}$ is
$\varphi_{i}^{{\text{\tiny LOD}}}=\mathcal{H}^{-1}\varphi_{i}^{H},\quad i=1,\dots,N_{H}$ (3.3)
where $\varphi^{H}_{i}\in\mathbb{P}^{1}(\mathcal{T}_{H})$ denote the standard hat function associated with the $i$th node of $\mathcal{T}_{H}$. But this basis is no longer locally supported, which, in contrast to its good approximation properties, gives the impression that $V_{\text{\tiny LOD}}$ is not a suitable approximation space, since it implies densely populated system matrices. However, it is known that the basis functions (3.3) decay exponentially (cf. [13, 14]) and can therefore be restricted to local patches by making small localization errors. We make this more precise in the following:
@@
and $\varphi_{i}^{\text{\tiny LOD}}=(I-\mathcal{C})\varphi^{H}_{i}$, $i=1,\dots N_{H}$. For $\ell\in\mathbb{N}$ we next define the $\ell$-patch of $K\in\mathcal{T}_{H}$ recursively by
@@
Now we define the local element corrector $\mathcal{C}^{\ell}_{K}:H^{1}_{0}(\mathcal{D})\rightarrow W^{\ell}_{K}$ of $K$ via
@@
For the implementation of the LOD one has to solve the local problems (3.4) for $v=\varphi^{H}_{i}$ which is done in practice by a fine scale discretization of mesh size $0<h<H$, see e.g. [7]. We choose a standard $\mathbb{P}^{1}$-Lagrange FEM on a fine triangulation $\mathcal{T}_{h}$ which we assume for simplicity to be a refinement of the coarse mesh $\mathcal{T}_{H}$. This leads us to fully-discrete spatial approximation space
@@ 4 Time Integration and Fully-Discrete Schemes
$\int_{0}^{T}(\mathrm{i}\partial_{t}u_{\mathbb{P}^{k}},v)-(u_{\mathbb{P}^{k}},v)_{\mathcal{H}}-\beta\big{(}(|u_{\mathbb{P}^{k}}|^{2})u_{\mathbb{P}^{k}},v\big{)}\,dt=0\quad\forall v\in\tilde{W}^{\ell,h}_{q-1,k}$ (4.3)
@@ 5 Numerical Results
For simplicity, we calculate the $H^{1}$-error via the quantity $\|u_{\mathrm{ref}}(T)-u_{{\text{\tiny LOD}}}(T)\|_{H^{1}}$ for the cG-LOD-method where $u_{\mathrm{ref}}$ is a reference solution obtained by the cG-$\mathbb{P}^{1}$-method on the fine mesh with mesh size $h=30\cdot 2^{-16}$. The error is calculated accordingly for the cG-$\mathbb{P}^{k}$-methods with the same reference solution. The CPU times shown in Fig. 5.2 and Fig. 5.4 are in any case measured for the time stepping (online), excluding pre-computations (offline) such as the calculation of the LOD basis functions. There are two reasons for this: First, the calculations of the basis functions are completely independent and can be calculated in parallel. This allows the computational costs to be kept small on modern computer architectures. The second reason is that, especially on long time scales, the costs for the time stepping clearly dominate the costs for the pre-computations and are therefore negligible.

### v2

@@ 3 Spatial Discretization by Localized Orthogonal Decomposition
A natural choice for a basis of $V_{\text{\tiny LOD}}$ is $\mathcal{H}^{-1}\varphi_{i}^{H}$, $i=1,\dots,N_{H}$ where $\varphi^{H}_{i}\in\mathbb{P}^{1}(\mathcal{T}_{H})$ denotes the standard hat function associated with the $i$th node of $\mathcal{T}_{H}$. But this basis (and in general any other basis) is no longer locally supported, which, in contrast to its good approximation properties, gives the impression that $V_{\text{\tiny LOD}}$ is not a suitable approximation space as it implies densely populated system matrices. However, it is known that $V_{\text{\tiny LOD}}$ is spanned by a basis whose functions decay exponentially (cf. [13, 14]) and can therefore be restricted to local patches by making small localization errors. We make this more precise in the following:
@@
and is spanned by the (exponentially decaying) basis $\varphi_{i}^{\text{\tiny LOD}}:=(I-\mathcal{C})\varphi^{H}_{i}$, $i=1,\dots N_{H}$. For $\ell\in\mathbb{N}$ we introduce the $\ell$-patch of an element $K\in\mathcal{T}_{H}$ recursively by
@@
Further, we define the local element corrector $\mathcal{C}^{\ell}_{K}:H^{1}_{0}(\mathcal{D})\rightarrow W^{\ell}_{K}$ of $K$ via
@@
For the implementation of the LOD one has to solve the local problems (3.3) for $v=\varphi^{H}_{i}$ which is done in practice by a fine scale discretization of mesh size $0<h<H$, see e.g. [7]. We choose a standard $\mathbb{P}^{1}$-Lagrange FEM on a fine triangulation $\mathcal{T}_{h}$ which we assume for simplicity to be a refinement of the coarse mesh $\mathcal{T}_{H}$. This leads us to the fully-discrete spatial approximation space
@@ 4 Time Integration and Fully-Discrete Schemes
$\int_{0}^{T}(\mathrm{i}\partial_{t}u_{\mathbb{P}^{k}},v)-(u_{\mathbb{P}^{k}},v)_{\mathcal{H}}-\beta\big{(}|u_{\mathbb{P}^{k}}|^{2}u_{\mathbb{P}^{k}},v\big{)}\,dt=0\quad\forall v\in\tilde{W}^{\ell,h}_{q-1,k}$ (4.3)
@@ 5 Numerical Results
For simplicity, we calculate the $H^{1}$-error via the quantity $\|u_{\mathrm{ref}}(T)-u_{{\text{\tiny LOD}}}(T)\|_{H^{1}}$ for the cG-LOD-method where $u_{\mathrm{ref}}$ is a reference solution obtained by the cG-$\mathbb{P}^{1}$-method on the fine mesh with mesh size $h=30\cdot 2^{-16}$. The error is calculated accordingly for the cG-$\mathbb{P}^{k}$-methods with the same reference solution. The CPU times shown in Fig. 5.2 and Fig. 5.4 are in any case measured for the time stepping (online), excluding pre-computations (offline) such as the calculation of the LOD basis functions. There are two reasons for this: First, the calculations of the basis functions are completely independent and can be calculated in parallel. This allows the computational costs to be kept small on modern computer architectures. The second reason is that, especially on long time scales, the costs for the time stepping clearly dominate and the costs for the pre-computations are therefore negligible.